Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: A/dual pol
Obs time (UT): EU 03:30(19/02)-14:30(19/02)
EU Mk3 Tapes/St.: 3
To give feedback on this experiment, click here
Effelsberg (dgraham at Wed Feb 19 09:41:44 EST 1997) Minor failures
Weather and ONSOURCE commands off until 1048. Otherwise OK
Medicina (Orfei at Wed Feb 19 09:43:27 EST 1997) Minor failures
Due to check2a1 and check2a2 duration time lost some minute
for some scans until 050121500. Footage doesn`t match with the
Torun (R.Feiler at Wed Feb 19 20:39:23 EST 1997) Success
We had to finish at 13:48 - sources outside elevation limits (3 deg.).
Westerbork (Tony Foley at Thu Feb 20 05:25:35 EST 1997) Minor failures
Problems with pointing of one telescope (RT8)
so it was removed from the array 0959UT to 1500UT
<P><DT><STRONG>Jodrell</STRONG> (Alastair Gunn at Wed Feb 26 09:17:42 EST 1997) <EM>Minor failures</EM>
POLARISATIONS ARE CROSSED. Telescope stowed at about 09:00 due
to high winds. Tapes JODR0037,7740 NRAOJPL1,6547 and MEDP1022,4248
sent to Bonn via Woodcroft on 24/02/97.
<P><DT><STRONG>Cambridge</STRONG> (Alastair Gunn at Wed Feb 26 09:21:12 EST 1997) <EM>Minor failures</EM>
Problems with wind - Telescope parked at about 09:25, back on-line
at about 11:00, offline again at about 13:00. Tapes MEDI3124,9E7D
USNO0350,B244 and IAA00234 sent to Bonn via Woodcroft on 24/02/97.
Calibration (Fredrik T Rantakyro at Thu May 8 11:32:43 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the
Feb97 directory with the name es16b.antab. This file contains the
preliminary calibration results and may be updated when I receive
updated information from the telescopes and the results from the
datareduction of the calibration experiments.
MPIfR (Scott Aaron at Tue Sep 23 14:42:48 METDST 1997) Minor failures
Torun tapes unreadable.
MPIfR (Scott Aaron at Thu Oct 2 14:31:37 METDST 1997) Success
Correction: Torun tapes readable, data in upper four IFs only.
MPIfR (Scott Aaron at Fri Mar 13 10:53:35 MET 1998) Minor failures
Correlation completed. Very bad tracking problems with Torun tapes;
correlation of those baselines not completed.